1/1

Михаил Лермонтов,  Михаил Лермонтов,
Ангел  The Angel

übersetzt von N. N. 97

По небу полуночи ангел летел, И тихую песню он пел, И месяц, и звезды, и тучи толпой Внимали той песне святой.

At midnight an angel was crossing the sky,

And quietly he sang;

The moon and the stars and the concourse of clouds Paid heed to his heavenly song.

Он пел о блаженстве безгрешных духов Под кущами райских садов, О Боге великом он пел, и хвала Его непритворна была.

He sang of the bliss of the innocent souls

In heavenly gardens above;

Of almighty God he sang out, and his praise

Was pure and sincere.

Он душу младую в объятиях нес Для мира печали и слез; И звук его песни в душе молодой Остался - без слов, но живой.

He bore in his arms a young soul To our valley of sorrow and tears; The young soul remembered the heavenly song So vivid and yet without words.

И долго на свете томилась она, Желанием чудным полна, И звуков небес заменить не могли Ей скучные песни земли.

And long did it struggle on earth, With wondrous desire imbued; But none of the tedious songs of our earth Could rival celestial song.

Verfasser dieses russischen Gedichtes ist Михаил Лермонтов (∗1814-10-15 – †1841-07-27).

Arg-1039-2110 (2014-06-03 12:23:12)

Pri la poeto Lermontov vidu la retejon http: // eo. wikipedia. org/ wiki/ MiД�ail_Lermontov .

Übersetzung des russischen Gedichtes “Ангел” von Михаил Лермонтов (∗1814-10-15 – †1841-07-27) ins Englische durch N. N. 97.

Arg-1039-2111 (2014-06-03 11:29:07)

Tiu ĉi laŭvorta poem-angligo troviĝas en http: // max. mmlc. northwestern. edu/ ~mdenner/ Demo/ texts/ angel. html .

http://www.poezio.net/version?poem-id=1039&version-id=,2110,2111